OpenStudy (anonymous):

Find the average rate of change of the function y = x^2 - 3x + 5 over the interval [-1,3]

OpenStudy (amistre64):

\[\frac{f(3)-f(-1)}{3--1}\]

OpenStudy (amistre64):

hopefully, you know what f(3) and f(-1) mean :)

OpenStudy (amistre64):

f(3) = 3^2 - 3(3) + 5 = 5 f(-1) = (-1)^2 -3(-1) + 5 = 10

OpenStudy (amistre64):

5 - 10 ----- = avg rate of change 4

OpenStudy (amistre64):

3+1 = 4 + 5 = 9 , not 10

OpenStudy (amistre64):

sooo close lol

OpenStudy (amistre64):

5-9 ---- = -1 then :) 4

OpenStudy (anonymous):

yes, thank you it is the derivative of b minus deriv of a, all over function b - function a

OpenStudy (anonymous):

right...

OpenStudy (amistre64):

derivatives are instantaneous rates of change; average rates are just last - first divided by well, b-a

OpenStudy (amistre64):

i wonder if derivs work out as well 2x-3; 2-3 = -1 , 6-3 = 3 -1-3 ---- = -1 hmmm, might be on to something there 3+1

myininaya (myininaya):

you are just looking for the slope of the secant line that lies on the curve y=x^2-3x+5 touching the two points (-1,f(-1)) and (3,f(3))

OpenStudy (amistre64):

nope, 2(-1)-3 = -5

myininaya (myininaya):

so basically you learn how to find average rate of change in algebra

myininaya (myininaya):

its just the slope of a line

OpenStudy (amistre64):

farm says I shouldnt harass you for your banning techniques :)

myininaya (myininaya):

shhh i will accidentlly ban you

OpenStudy (amistre64):

:) practice makes perfect

OpenStudy (anonymous):

wait, i am confused a little bit, what do i do in simple steps to solve this?

OpenStudy (amistre64):

(3^2-3(3)+5) - ((-1)^2-3(-1)+5) ----------------------------- 3-(-1)

OpenStudy (amistre64):



OpenStudy (anonymous):

that is equal to -1 right

OpenStudy (amistre64):

the avg rate of change is just the slope of the line between 2 points

OpenStudy (amistre64):

the instant rate of change is the slope of the tangent line to the curve at a point; also known as the derivative

OpenStudy (anonymous):

okay, i'm good

OpenStudy (amistre64):

yay!! :)